OpenStudy (anonymous):

Given that f(x) = 2x + 1 and g(x) = the quantity of 3x minus 1, divided by 2, solve for g(f(3)). 5 7 9 10

Vocaloid (vocaloid):

well, start from the inside and work your way out f(3) = ?

OpenStudy (08surya):

10

OpenStudy (08surya):

g(x)=(3x-1)/2,f(x)=2x+1,g(f(3))=? f(3)=2*3+1=7 g(7)=(3*7-1)/2=10 ans.